OpenStudy (anonymous):

Let f(x)=(x−6)2 . Find the average rate of change of f(x) with respect to x from x=a to x=a+h if h≠0

myininaya (myininaya):

\[\frac{f(x_1)-f(x_0)}{x_1-x_0}\] Just use the formula for slope

OpenStudy (anonymous):

ummm how would that end up giving me the right answer? isnt there more to the problem?

OpenStudy (anonymous):

im still just really confused haha

myininaya (myininaya):

you are given x_1=a and x_0=a+h

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

can you show me the step by step way on how to do this please? It would help me a lot :) I'm sorry im not not fully understanding it

myininaya (myininaya):

we plug in

myininaya (myininaya):

\[\frac{f(a)-f(a+h)}{a-(a+h)}\]

OpenStudy (anonymous):

so is that the final answer?

myininaya (myininaya):

no we have to replace f(a) with (a-6)^2 and f(a+h) with (a+h-6)^2

myininaya (myininaya):

Then simplify using algebra

myininaya (myininaya):

You will probably have to multiply some stuff out

myininaya (myininaya):

\[\frac{(a-6)^2+(a+h-6)^2}{a-(a+h)}\] Can you play with this and simplify?

myininaya (myininaya):

oops I type-o \[\frac{(a-6)^2-(a+h-6)^2}{a-(a+h)}\]

myininaya (myininaya):

I will show you a trick that might be helpful: \[(a+h-6)^2=(a+h-6)(a+h-6)=([a-6]+[h])([a-6]+[h]) \] \[=(a-6)^2+2(a-6)h+h^2 =a^2-12a+36+2a-12+h^2\]

OpenStudy (anonymous):

Thank you!